OpenStudy (kaydra14):

How many equal sections divide a directed line segment if it is to be partitioned with a 9:129:12 ratio?

OpenStudy (518nad):

okay first can we simplify this ratio down?

OpenStudy (518nad):

hey?

OpenStudy (518nad):

are you there?

OpenStudy (kaydra14):

yes

OpenStudy (518nad):

okay lets start with simplifying this ratio down

OpenStudy (518nad):

3 is common in all 3 numbers factor 3 out

OpenStudy (518nad):

3*3:3*43:3*4 =3:43:4

OpenStudy (mathmale):

In this problem we're partioning (dividing up) a directed line segment in the ratio 9:129:12. Each partition is to have the same length. You cannot tell one partiion from another. Think: 9 partitions plus 129 more, identical partitions, plus 12 more, identical partitions. How many partitions will there be, altogether? Now suppose the question were different: "What fraction of the total number of partitions constitutes the first 9 partitions? You'd add up the counts representing the 3 partitions and divide that 9 by that sum. What is the sum? What is the ratio 9:sum? Here's where Dan's suggestion makes sense. 9/sum would be the same ratio AFTER factoring out the common ratio, 3. Experiment with this. Try to answer the original question first. If you're still unsure, ask a clarifying question or two.

OpenStudy (sshayer):

\[\frac{ 9 }{ 129 }=\frac{ 12 }{ x },9x=129 \times 12,x=?\]